OpenStudy (anonymous):

Simplify the expression 6x^2(3x)

OpenStudy (anonymous):

\[6x^2(3x)=6×3×x^2×x=18x^3\]

OpenStudy (anonymous):

6x^2(3x) = 18 x^3